OpenStudy (anonymous):

put (0,3) (4,0) in the form y = mx + b, where b is the constant.

OpenStudy (saifoo.khan):

y = mx + b Where m is the slope and b is the y-intercept.

OpenStudy (saifoo.khan):

Step 1, find the slope. \[m = \frac{y_2- y_2}{x_2- x_1} \to \frac{0 -3}{4-0} \to \frac{-3}{4}\]

OpenStudy (saifoo.khan):

@Crissssssssssssss , please use the "attach file" option to upload a file.

OpenStudy (saifoo.khan):

Ok, cool. next, since you already know that y-intercept is 3, so you will rewrite the equation as, y = -3/4 x + 3

OpenStudy (anonymous):

Is that all?

OpenStudy (saifoo.khan):

Yes.